### v1

Abstract
Missing modalities pose a major issue in Alzheimer’s Disease (AD) diagnosis, as many subjects lack full imaging data due to cost and clinical constraints. While multi-modal learning leverages complementary information, most existing methods train only on complete data, ignoring the large proportion of incomplete samples in real-world datasets like ADNI. This reduces the effective training set and limits the full use of valuable medical data. While some methods incorporate incomplete samples, they fail to effectively address inter-modal feature alignment and knowledge transfer challenges under high missing rates. To address this, we propose a Prototype-Guided Adaptive Distillation (PGAD) framework that directly incorporates incomplete multi-modal data into training. PGAD enhances missing modality representations through prototype matching and balances learning with a dynamic sampling strategy. We validate PGAD on the ADNI dataset with varying missing rates (20%, 50%, and 70%) and demonstrate that it significantly outperforms state-of-the-art approaches. Ablation studies confirm the effectiveness of prototype matching and adaptive sampling, highlighting the potential of our framework for robust and scalable AD diagnosis in real-world clinical settings.
Keywords:
Alzheimer’s Disease, Missing Modalities, Incomplete Data Training, Prototype Matching, Adaptive Sampling.
1 Introduction
Alzheimer’s Disease (AD) is a common neurodegenerative disorder that leads to irreversible cognitive decline, posing major challenges to global healthcare, especially as the population ages [1, 2]. Early diagnosis, particularly at the Mild Cognitive Impairment (MCI) stage, is crucial for slowing disease progression [3, 4, 5]. Multi-modal neuroimaging, including Magnetic Resonance Imaging (MRI) and Positron Emission Tomography (PET), is widely used in AD research due to its ability to capture complementary structural and functional brain changes [6, 7, 25]. However, PET scans are expensive and not always accessible, leading to a significant portion of subjects in existing datasets lacking PET imaging. This reduces the availability of complete multi-modal data for AD studies [8]. As a result, many models cannot utilize data instances missing this modality, limiting their ability to fully exploit cross-modal information.
Most studies addressing this issue train on complete multi-modal data and later adapt to single-modality inference using techniques like knowledge distillation. This assumes full modality availability during training, making it less effective for real-world scenarios where missing data is common. Only a few works [26] have attempted to train directly on incomplete multi-modal data. However, without incorporating incomplete data into the teacher network’s training process, issues like unreliable inter-modal feature alignment [9, 11] and inefficient knowledge transfer [10, 12] become more pronounced, reducing generalizability and diagnostic reliability.
To address this, we propose the PGAD framework, which integrates two key components: Prototype Consistency Matching (PCM) and Adaptive Multi-Modal Sampling (AMS). PCM reduces feature misalignment by learning class-level prototypes from complete multi-modal samples and enforcing consistency between missing modality features and their corresponding prototypes. This helps unpaired MRI samples retain essential modality-specific characteristics, minimizing feature discrepancies between missing and available modalities. Meanwhile, AMS improves the transfer of fused multi-modal features to MRI representations, ensuring that useful cross-modal information is effectively leveraged despite missing data.
In summary, our contributions are threefold: (1) We introduce PGAD, a novel knowledge distillation framework that directly incorporates incomplete multi-modal data into training, enhancing feature learning under missing conditions. (2) We propose PCM and AMS, which effectively mitigate modality gaps by aligning feature representations and dynamically balancing data utilization, improving model robustness. (3) We conduct extensive experiments on the ADNI dataset, demonstrating that PGAD significantly outperforms existing approaches, achieving state-of-the-art performance in both AD classification and MCI conversion prediction.
2 Method
This study focuses on improving AD diagnosis under missing modality conditions by proposing the PGAD framework. PGAD is designed to enhance feature alignment, facilitate effective knowledge transfer, and stabilize training in the presence of incomplete multi-modal data. It achieves this by leveraging class prototypes to guide unpaired samples and employing a dynamic sampling strategy to regulate modality distributions. These mechanisms collectively improve diagnostic accuracy and model robustness. The overall framework is illustrated in Fig. 1.
Figure 1: Overview of the PGAD framework. The teacher network utilizes both MRI and PET data to extract joint feature representations and construct category prototypes. The student network, trained with only MRI data, operates under two conditions: (1) paired MRI samples ($H_{pair}^{mri}$) with corresponding PET data are constrained using distillation loss, and (2) unpaired MRI samples ($H_{unpair}^{mri}$) rely on category prototypes via the PCM module. The AMS module regulates the balance of paired and unpaired samples, enhancing training stability and cross-modal feature alignment.
2.1 PGAD Framework Overview
The PGAD framework is designed to bridge the gap between complete and incomplete multi-modal learning by leveraging prototype-based feature alignment and adaptive sampling strategies. PGAD introduces two key components: PCM and AMS. PCM aligns unpaired MRI samples with category prototypes extracted from MRI-PET paired data, ensuring missing modality features retain meaningful representations. AMS dynamically adjusts the balance of paired and unpaired samples, mitigating data imbalance issues and improving cross-modal feature learning.
By combining prototype-guided knowledge transfer and structured sampling strategies, PGAD enables the student model to utilize available multi-modal information effectively while maintaining robustness when faced with missing modalities. This approach enhances feature consistency, stabilizes training, and improves generalization in real-world missing modality scenarios.
2.2 Prototype Consistency Matching (PCM)
Prototype Consistency Matching (PCM) aims to address feature misalignment caused by missing modalities by enforcing consistency between unpaired MRI samples and class-specific prototypes. Given a set of complete multi-modal samples, PCM constructs category prototypes by averaging the feature embeddings of MRI-PET paired data:
$z_{c}=\frac{1}{N_{c}}\sum_{i\in C_{c}}h_{i}^{all},$ (1)
where $h_{i}^{all}$ represents the joint MRI-PET feature embedding for a paired sample, and $N_{c}$ is the number of samples in class $c$. For an unpaired MRI sample, the closest prototype is assigned, and feature alignment is enforced through a consistency loss:
$L_{\text{proto}}=\frac{1}{|\mathcal{U}|}\sum_{i\in\mathcal{U}}\|h_{i}^{mri}-z_{c}^{*}\|^{2}$ (2)
where $h_{i}^{mri}$ is the MRI feature embedding of an unpaired sample, and $z_{c}^{*}$ is the nearest class prototype. This constraint ensures that missing modality samples retain class-specific information, improving representation learning under incomplete conditions.
2.3 Adaptive Multi-Modal Sampling (AMS)
Adaptive Multi-Modal Sampling (AMS) dynamically adjusts the proportion of paired and unpaired samples in each training batch to ensure stable learning across different missing rates. Instead of relying on a fixed sampling ratio, AMS introduces an adaptive weighting mechanism:
$w=\frac{1}{1+e^{-\alpha(r-\tau)}},$ (3)
where $r$ is the current missing rate, $\tau$ is a threshold parameter, and $\alpha$ controls the sharpness of the transition. This formulation allows the model to prioritize paired samples when missing rates are low while effectively learning from unpaired samples under higher missing rates.
To further enhance robustness, AMS employs a contrastive objective to encourage discriminative feature learning between true paired samples and randomly assigned unpaired samples:
$L_{\text{pair}}=-\frac{1}{|\mathcal{P}|}\sum_{(i,j)\in\mathcal{P}}\log\frac{\exp(\text{sim}(h_{i}^{mri},h_{j}^{pet})/\tau)}{\sum_{k}\exp(\text{sim}(h_{i}^{mri},h_{k}^{pet})/\tau)},$ (4)
where $\text{sim}(\cdot,\cdot)$ measures feature similarity, and $\tau$ is a temperature parameter. This strategy ensures that unpaired samples learn modality-invariant features, improving cross-modal feature alignment.
2.4 Optimization Strategy
The Prototype-Guided Adaptive Distillation (PGAD) framework is optimized through a combination of five loss terms that jointly enhance feature learning, knowledge transfer, and cross-modal alignment. The overall objective function is formulated as:
$L=\lambda_{\text{tea}}L_{\text{tea}}+\lambda_{\text{stu}}L_{\text{stu}}+\lambda_{\text{kl}}L_{\text{kl}}+\lambda_{\text{pair}}L_{\text{pair}}+\lambda_{\text{proto}}L_{\text{proto}}.$ (5)
The teacher classification loss $L_{\text{tea}}$ supervises the multi-modal teacher network to ensure effective feature extraction, while the student classification loss $L_{\text{stu}}$ guides the single-modality student model, enforcing consistency between paired and unpaired samples. The knowledge distillation loss $L_{\text{kl}}$ aligns the output distributions of the teacher and student models, enabling effective transfer of multi-modal knowledge. To enhance feature alignment, the paired sample consistency loss $L_{\text{pair}}$ constrains representations extracted from paired MRI and PET inputs, ensuring they retain modality-specific but complementary information. Finally, the prototype consistency loss $L_{\text{proto}}$ mitigates modality gaps by aligning unpaired MRI samples with category prototypes, preserving class-discriminative features and improving feature robustness. By integrating these five loss components, PGAD effectively balances classification, knowledge transfer, and feature consistency, ensuring stable learning under varying missing modality conditions and achieving superior performance in AD classification and MCI conversion prediction.
3 Experiments
3.1 Data Acquisition and Processing
3.1.1 Dataset:
This study utilized data from the Alzheimer’s Disease Neuroimaging Initiative (ADNI), including ADNI-1, ADNI-2, ADNI-GO, and ADNI-3. The selected samples all contain both T1-weighted MRI and FDG-PET modalities to ensure data completeness and consistency. The final dataset consists of 394 Alzheimer’s Disease patients, 433 CN individuals, 185 pMCI patients, and 245 sMCI patients. The pMCI group includes individuals initially diagnosed with MCI who progressed to AD within 36 months, while the sMCI group comprises individuals who remained clinically stable as MCI for at least 36 months. Since specific age and gender information is not directly provided in the dataset, we estimated these attributes based on typical distributions in the ADNI database.
Table 1: Demographic information of the dataset
Class Number Gender (Male/Female) Age (Mean ± SD)
AD 394 45% / 55% 74 ± 8
CN 433 47% / 53% 72 ± 7
pMCI 185 46% / 54% 73 ± 6
sMCI 245 48% / 52% 71 ± 7
3.1.2 Preprocessing:
MRI preprocessing was performed using FSL, including reorientation, field of view estimation, skull stripping, and affine registration to the MNI152 template. PET preprocessing was conducted with SPM12, involving realignment, normalization, coregistration with MRI, image calculation, and smoothing. This ensures both modalities are spatially aligned for further analysis.
3.1.3 Experimental Settings:
We evaluate the PGAD framework on Alzheimer’s Disease classification (AD vs. CN) and MCI conversion prediction (pMCI vs. sMCI). The dataset is split using stratified five-fold cross-validation, ensuring a balanced distribution of classes across folds. To simulate missing modalities, experiments are conducted under incomplete rates of 20%, 50%, and 70%. During training, the teacher network processes both paired (MRI+PET) and unpaired (MRI with randomly assigned PET) data, while the student network is trained using only MRI. Final results are reported as the average performance across all five folds to ensure robustness and reliability.
The model is trained on four NVIDIA RTX 4090 GPUs with a batch size of 32 using the Adam optimizer ($1\times 10^{-4}$ initial learning rate, cosine annealing, weight decay $5\times 10^{-5}$) for 100 epochs. The loss function consists of multiple terms with weights: $\lambda_{\text{tea}}=1.0$, $\lambda_{\text{stu}}=1.0$, $\lambda_{\text{kl}}=0.5$, $\lambda_{\text{pair}}=0.5$, and $\lambda_{\text{proto}}=0.5$.
3.2 Comparison with Existing Approaches
We compare PGAD with existing methods for Alzheimer’s disease classification under AD vs. CN and pMCI vs. sMCI tasks. As shown in Table 2, PGAD consistently achieves superior performance across all metrics.
For fairness, PGAD is trained on complete multi-modal data and evaluated with MRI-only inputs, aligning with prior knowledge distillation-based methods. The results of other approaches are taken from their respective publications.
In AD vs. CN classification, PGAD outperforms MRI-only models and existing missing modality adaptation methods, demonstrating better feature alignment and robustness. It improves sensitivity while maintaining high specificity, ensuring balanced classification. For pMCI vs. sMCI, PGAD achieves higher accuracy and stability across missing modality conditions. It effectively distinguishes progressive MCI cases, highlighting its potential for early risk identification.
These results confirm PGAD’s effectiveness in handling missing modality learning, offering significant improvements over MRI-only and adaptation-based approaches. By leveraging prototype alignment and adaptive sampling, PGAD ensures high diagnostic accuracy and robustness, making it suitable for real-world clinical applications.
Table 2: Comparisons of classification performance across different methods on AD vs. CN and pMCI vs. sMCI tasks. The best and second-best results in each metric are highlighted in bold and underlined.
Task Method Metrics
AUC (%) ACC (%) SEN (%) SPE (%)
AD vs. CN 3D-Mixer [23] 96.8 90.6 - -
DA-MIDL [22] 96.5 92.4 91.0 93.8
GF-NET [24] 93.8 93.2 94.6 94.2
DFTD [11] 96.9 84.2 91.7 93.7
Wang et al. [25] 93.3 87.5 87.4 87.5
MDT-Student et al. [26] 79.6 79.4 68.0 78.6
PGAD (Ours) 97.2 93.5 94.0 94.5
pMCI vs. sMCI 3D-Mixer [23] 77.6 75.5 - -
DA-MIDL [22] 85.2 80.2 77.1 82.6
DFTD [11] 83.8 51.2 73.0 79.0
PGAD (Ours) 85.5 80.6 77.6 81.4
3.3 Ablation Study
The ablation study in Tab. 3 highlights the contribution of PCM and AMS in handling missing modalities. Without PET, classification performance declines significantly, emphasizing the importance of multi-modal information. However, incorporating PCM and AMS mitigates this degradation. PCM improves feature alignment by learning class-level prototypes, ensuring missing modality features retain class-discriminative information. AMS dynamically balances paired and unpaired samples, preventing overfitting and enhancing cross-modal knowledge transfer. Our method demonstrates strong robustness, maintaining stable performance across different missing rates. While performance naturally declines as missing data increases, PCM preserves structured feature representations, and AMS prevents training imbalance, ensuring reliable classification even under severe modality loss.
Table 3: Ablation study on the effect of PCM and AMS under different missing modality ratios.
Missing Ratio PCM AMS ACC (%) SEN (%) SPE (%) AUC (%)
Full PET 93.5 94.0 94.5 97.2
0.2 $\times$ $\times$ 91.12 96.42 90.72 94.50
$\times$ 92.35 96.42 91.74 94.94
92.74 97.61 92.68 95.98
0.5 $\times$ $\times$ 91.12 97.61 90.47 93.65
$\times$ 91.93 92.85 95.40 95.06
91.93 95.23 92.53 94.75
0.7 $\times$ $\times$ 87.90 95.23 86.32 89.93
$\times$ 88.70 95.23 87.41 91.10
89.51 92.85 90.86 92.53
3.4 Visualization Analysis of Experimental Results
To further evaluate the effectiveness of our model in identifying Alzheimer’s disease-related biomarkers, we conducted a visualization analysis of its attention regions, as shown in Fig. 2. The figure overlays Grad-CAM heatmaps on MRI axial, coronal, and sagittal slices to highlight the model’s primary focus areas.
The results show that the model consistently attends to key brain regions associated with Alzheimer’s pathology, including the hippocampus, lateral ventricles, and cortical atrophy areas. These highlighted regions align with findings from medical research and have been confirmed through discussions with clinical experts, indicating that the model effectively captures disease-relevant features.
Moreover, the attention maps remain stable across different subjects, demonstrating the model’s robustness in feature localization. This further validates its generalizability in Alzheimer’s classification and highlights its potential for real-world clinical applications.
Figure 2: Visualization of the model’s attention regions using saliency map overlaid on MRI slices. The red-highlighted areas indicate the regions most relevant to Alzheimer’s disease classification in axial, coronal, and sagittal views.
4 Conclusion
We introduced the PGAD framework to address the challenges of aligning missing modality features and improving knowledge transfer in Alzheimer’s Disease diagnosis. By integrating Prototype Consistency Matching (PCM) and Adaptive Multi-Modal Sampling (AMS), PGAD effectively leverages incomplete multi-modal data, mitigating feature misalignment and enhancing cross-modal learning. Experimental results and ablation studies validate its effectiveness over existing methods.These findings highlight the potential of prototype-guided learning and adaptive sampling for improving multi-modal medical imaging under real-world missing data conditions.